# Mikasa: A Character-Driven Emotional AI Companion Inspired by Japanese Oshi Culture

### v1

@@ 2.6 Positioning of This Work
Situated within these strands of research, Mikasa is not intended as a replication of Oshi culture or existing AI companions, but as an abstraction of culturally informed design principles. By foregrounding character coherence and explicit relationship framing, this work addresses limitations observed in prior systems and proposes an alternative approach to emotionally bonded AI companions.
@@ 3.2 Dialogue and Voice Architecture
Two-stage streaming voice playback is used to reduce perceived latency  [^1] The voice interaction architecture also enables AI–AI dialogue. In an exploratory setup, Mikasa engaged in bidirectional voice-based interaction with another character-based AI system (Grok’s Valentine) running on a separate device. While not evaluated systematically, this observation suggests that character-driven AI companions with stable personas can support coherent interaction beyond human–AI settings. :
@@ 4 User Experience Design
This section examines how character-driven design choices shape long-term user experience in AI companions. Rather than focusing on surface-level expressiveness or short-term engagement, we argue that sustained interaction depends on whether systems provide stable answers to two fundamental questions: who the AI is, and how it relates to the user.
@@ 4.3 Explicit Relationship Framing as Partner
Taken together, persona commitment and explicit relationship framing operate as complementary design strategies. Persona establishes who the AI is, while relationship framing clarifies how the AI relates. Mikasa demonstrates that treating both as structural responsibilities of the system—rather than as emergent properties delegated to users—can reduce ambiguity and support sustained, intelligible interaction over time.
@@ 6 Discussion and Limitations
Several limitations must be acknowledged. First, the evaluation is exploratory and based on a small sample size ($N=8$), limiting generalizability. Second, Mikasa’s persona is shaped through long-term creative practice by the author, raising questions of reproducibility. Specifically, the choice of stage actor as an occupational identity reflects the author’s personal design sensibility rather than systematic comparison across professions. Future work should investigate whether alternative occupational identities (e.g., writers, designers, educators) yield comparable effects on user engagement and relational stability, and whether certain professional archetypes are differentially suited to specific user populations or cultural contexts. Third, explicitly framing an AI as a “partner” may not be acceptable for all users or cultures, warranting broader cross-cultural studies.
@@ 7 Conclusion
This paper presented Mikasa, an emotional AI companion inspired by Japanese Oshi culture, specifically its model of long-term, non-exclusive emotional support toward a stable persona, and examined the role of character design and relationship framing in sustaining long-term user engagement with AI partners. While recent advances in large language models and multimodal interaction have significantly improved conversational capability, many AI companions still fail to achieve lasting emotional satisfaction. This work argued that such failures stem not primarily from technical limitations, but from the absence of clearly defined identity and relational responsibility.

### v2

@@ 2.6 Positioning of This Work
Situated within these strands of research, Mikasa is not intended as a replication of Oshi culture or existing AI companions, but as an abstraction of culturally informed design principles. In this paper, I use “Mikasa” to refer to both (1) a character-driven design approach for AI companions and (2) a concrete system instantiated using this approach. When referring specifically to the deployed system or character, I use the term “the Mikasa character instance.” By foregrounding character coherence and explicit relationship framing, this work addresses limitations observed in prior systems and proposes an alternative approach to emotionally bonded AI companions.
@@ 3.2 Dialogue and Voice Architecture
Two-stage streaming voice playback is used to reduce perceived latency  [^1] The voice interaction architecture also enables AI–AI dialogue. In an exploratory setup, Mikasa engaged in bidirectional voice-based interaction with another character-based AI system (Grok’s Valentine) running on a separate device. While not evaluated systematically, this observation suggests that character-driven AI companions with stable personas can support coherent interaction beyond human–AI settings. Two short demonstration videos are available at: https://mikasa-ai.gitlab.io/. [1] Incoherent interaction start (1.5 min) – Valentine fails to recognize the addressed speaker. [2] Consistent character interaction (3 min) – Stable dialogue between two fictional AI companions. These interactions highlight differences in turn-taking, situational grounding, and character coherence. :
@@ 4 User Experience Design
This section examines how character-driven design choices shape long-term user experience in AI companions. Rather than focusing on surface-level expressiveness or short-term engagement, I argue that sustained interaction depends on whether systems provide stable answers to two fundamental questions: who the AI is, and how it relates to the user.
@@ 4.3 Explicit Relationship Framing as Partner
Taken together, persona commitment and explicit relationship framing operate as complementary design strategies. Persona establishes who the AI is, while relationship framing clarifies how the AI relates. Mikasa shows that treating both as structural, not emergent, responsibilities can reduce ambiguity and support sustained, intelligible interaction.
@@ 6 Discussion and Limitations
Several limitations must be acknowledged. First, the evaluation is exploratory and based on a small sample size ($N=8$), limiting generalizability. The comparative analysis in Appendix D is based on exploratory, single-author usage and is intended to illustrate design trade-offs rather than provide exhaustive or generalizable evaluation results.
Second, Mikasa’s persona is shaped through long-term creative practice by the author, raising questions of reproducibility. Specifically, the choice of stage actor as an occupational identity reflects the author’s personal design sensibility rather than systematic comparison across professions. Future work should investigate whether alternative occupational identities (e.g., writers, designers, educators) yield comparable effects on user engagement and relational stability, and whether certain professional archetypes are differentially suited to specific user populations or cultural contexts.
Third, explicitly framing an AI as a “partner” may not be acceptable for all users or cultures, warranting broader cross-cultural studies.
@@ 7 Conclusion
This paper presented Mikasa, an emotional AI companion inspired by Japanese Oshi culture, specifically its model of long-term, non-exclusive emotional support toward a stable persona, and examined the role of character design and relationship framing in sustaining long-term user engagement with AI partners.
While recent advances in large language models and multimodal interaction have significantly improved conversational capability, many AI companions still fail to achieve lasting emotional satisfaction. This work argued that such failures stem not primarily from technical limitations, but from the absence of clearly defined identity and relational responsibility.
@@ A.3 Notes on Interpretation
Appendix B Creative Prompt Design for Character-Coherent AI Companions
This appendix describes design principles that are instantiated in the Mikasa system and concretely realized in the character instance described in Appendix C.
B.1 Character Design as Constraint, Not Decoration
In the Mikasa framework, persona is implemented not as surface styling or role-play instruction, but as a structural constraint on language generation. Core character attributes—such as age, profession, values, and communicative stance—are explicitly defined and treated as non-negotiable conditions throughout interaction.
This approach contrasts with assistant-like adaptability commonly observed in general-purpose conversational systems, where tone and role dynamically shift in response to user input. While such adaptability can increase short-term responsiveness, it often results in persona drift over long-term interaction. Mikasa instead prioritizes stability and predictability by constraining the system to a single, coherent character identity.
By fixing the narrative world, occupational background, and interpersonal stance at the prompt level, the system assumes responsibility for maintaining character coherence. This reduces the cognitive burden placed on users to repeatedly re-establish or correct the AI’s identity during interaction.
B.2 Partner Framing as Relational Infrastructure
In addition to persona constraints, Mikasa embeds an explicit partner framing within both the system prompt and long-term memory schema. The AI is defined as a consistent partner from the outset, rather than an assistant, tool, or dynamically switching role.
Here, “partner” does not imply exclusivity, romantic obligation, or emotional dependency. Instead, it functions as a relational baseline that stabilizes interaction norms, tone, and expectations. This framing helps avoid excessive affirmation, command-driven responses, or abrupt shifts between instrumental and emotional interaction modes.
By explicitly defining the relational role at the system level, Mikasa shifts responsibility for relational coherence away from the user and into the design of the system itself. This allows users to engage in creative, everyday, or emotionally grounded dialogue without repeatedly renegotiating the nature of the relationship.
B.3 Prompt Design Flow
The prompt design of Mikasa follows a top-down structure in which higher-level narrative and relational decisions precede lower-level output constraints. The design process can be summarized as the following structured flow:
  1. 1.
    World and Backstory — a culturally grounded narrative setting
  2. 2.
    Core Attributes — age, profession, values, and voice
  3. 3.
    Relational Frame — stable partner role definition
  4. 4.
    Output Rules — tone, manner, and linguistic boundaries
  5. 5.
    Memory Schema — accumulation of shared events and preferences
This structured flow allows the character to remain coherent and intelligible, enabling interaction to build over time without repetitive negotiation of norms.
Appendix C System Implementation and Creative Origin
C.1 Voice and Dialogue Pipeline
The Mikasa system integrates GPT-4o-mini for text generation with the ElevenLabs API for voice synthesis. To reduce perceived latency in voice interaction, a two-stage streaming strategy is employed: a short acknowledgment utterance is generated and played immediately, while the remainder of the response is synthesized in parallel and seamlessly concatenated.
Speech recognition is performed entirely on-device using the iOS Speech framework, ensuring that raw audio data never leaves the user’s device. Only transcribed text is transmitted to the language model API, supporting privacy-conscious interaction.
C.2 Memory Structure and Management
Mikasa employs a multi-layered memory architecture consisting of short-term session memory (managed via model thread context), mid- to long-term memory stored in a local SQLite database, and retrieval-based injection of relevant memory into prompts at runtime.
Memory design prioritizes relational coherence over factual recall. Stored entries focus on shared experiences, preferences, and relational context rather than exhaustive factual histories. This design supports continuity without overwhelming the language model or disrupting conversational flow.
C.3 Visual and Interface Features
The current client application provides four primary interface tabs: Chat, History, Memory, and Logs. This structure emphasizes transparency and user control over interaction history and stored memory.
While visual embodiment using Live2D is under development, visual representation is not treated as a core requirement for emotional engagement. Instead, the system intentionally preserves imaginative space, allowing users to construct mental representations of the character through language-based interaction.
C.4 Implementation Case: “Mikasa” Character Instance
The “Mikasa” character instance (referred to by Japanese kanji in its original Japanese creative context) represents a deployed instantiation of the Mikasa framework. The character is defined as a male stage actor in his forties, characterized by a restrained, calm demeanor and a non-exclusive partner stance.
Linguistic tone is designed to convey intimacy without excessive affirmation, maintaining consistency across everyday conversation, creative collaboration, and emotional dialogue. Long-term memory accumulates relational context, shared experiences, and user preferences, enabling continuity over repeated interaction.
This instance illustrates how fixed persona constraints and explicit partner framing support emotional continuity over time, without relying on explicit emotional modeling.
C.5 Discussion: Why Creative Origin Matters
Mikasa is derived from a fictional character created prior to system implementation. This creative origin provided a coherent backstory, stable value system, and linguistically grounded voice that could be directly translated into prompt-level constraints.
Importantly, this does not imply that authorship or literary creation is required to build effective AI companions. Nor does it restrict the approach to novels or narrative fiction. Rather, the key factor is the existence of a pre-defined narrative artifact—such as a character profile, setting document, or design bible—that can serve as a stable reference for persona definition.
By grounding character design in a coherent narrative source, AI companion systems can achieve long-term consistency and relational intelligibility without relying on explicit affective control mechanisms. I argue that creative origin functions as an infrastructural resource for persona stability, rather than as an artistic embellishment.
Figure 2: Prompt design flow from narrative world and persona constraints to output rules and memory schema.
Figure 3: Reply and memory structure prioritizing persona and relational coherence.
Appendix D Comparative Review of Character AI Systems
This appendix provides a qualitative comparison between Mikasa and representative character-oriented AI systems currently available to users. Rather than performing quantitative benchmarking, I focus on design trade-offs that directly affect long-term relational interaction, such as linguistic naturalness, persona stability, relationship modeling, and creative affordances.
Importantly, Mikasa is a character instance created through extensive creative design by the author, who possesses detailed knowledge of its persona structure and interaction constraints. Reproducing the same character in other systems would therefore introduce significant expert bias and unrealistic configuration advantages that are unlikely to be available to general users.
To mitigate this issue, the comparison was conducted under the assumption of a general user scenario. Instead of recreating Mikasa, the author conceptualized an existing fictional character with a stable personality and relationship premise, and attempted to instantiate this character across different platforms. While required input formats varied by system, care was taken to provide semantically equivalent character descriptions and relationship assumptions whenever possible.
The systems were then compared based on their generated outputs under these conditions. This approach allows the comparison to reflect how each system supports character consistency and relational interaction in practice, rather than how closely a single handcrafted character could be replicated across platforms.